### v2

# Prior-Guided Diffusion Planning for Offline Reinforcement Learning

Abstract
Diffusion models have recently gained prominence in offline reinforcement learning due to their ability to effectively learn high-performing, generalizable policies from static datasets. Diffusion-based planners facilitate long-horizon decision-making by generating high-quality trajectories through iterative denoising, guided by return-maximizing objectives. However, existing guided sampling strategies such as Classifier Guidance, Classifier-Free Guidance, and Monte Carlo Sample Selection either produce suboptimal multi-modal actions, struggle with distributional drift, or incur prohibitive inference-time costs. To address these challenges, we propose Prior Guidance (PG), a novel guided sampling framework that replaces the standard Gaussian prior of a behavior-cloned diffusion model with a learnable distribution, optimized via a behavior-regularized objective. PG directly generates high-value trajectories without costly reward optimization of the diffusion model itself, and eliminates the need to sample multiple candidates at inference for sample selection. We present an efficient training strategy that applies behavior regularization in latent space, and empirically demonstrate that PG outperforms state-of-the-art diffusion policies and planners across diverse long-horizon offline RL benchmarks. Our code is available at https://github.com/ku-dmlab/PG.
1 Introduction
Offline reinforcement learning (RL) allows agents to learn effective policies from fixed, pre-collected datasets, eliminating additional environment interaction. This paradigm is especially beneficial in scenarios where exploration is prohibitively expensive, unsafe, or impractical (levine2020offline, ). A core challenge in offline RL arises from distributional shifts, which can lead learned policies to favor out-of-distribution (OOD) actions due to overestimated value functions. To mitigate this issue, many existing methods incorporate behavior regularization to keep the learned policy close to the behavior policy that generated the data (levine2020offline, ; wu2019behavior, ; fujimoto2021minimalist, ; xu2023offline, ).
Recently, diffusion-based planners have been proposed as a promising direction for offline RL. Leveraging their strengths in modeling long-range temporal dependencies and conditioning on auxiliary information (ho2020denoising, ; dhariwal2021diffusion, ; austin2021structured, ; li2022diffusion, ; ho2022video, ; khachatryan2023text2video, ), diffusion models are well-suited for sequence modeling in complex decision-making tasks. By conditioning on future return surrogates, diffusion models can generate trajectories biased toward high-value behaviors, thereby demonstrating superior performance over prior approaches in long-horizon navigation and manipulation benchmarks across diverse offline RL settings (janner2022planning, ; ajay2022conditional, ; liang2023adaptdiffuser, ; li2023hierarchical, ; chen2024simple, ).
Diffusion planners have employed various guided sampling strategies to steer trajectory generation toward high-return outcomes. Classifier Guidance (CG) (dhariwal2021diffusion, ) leverages a learned critic to directly guide the denoising process toward high-value trajectories. Classifier-Free Guidance (CFG) (ho2022classifier, ), in contrast, achieves conditioning by interpolating between conditional and unconditional diffusion models, removing the need for an additional network. While conceptually appealing, these methods face practical limitations: CG can produce suboptimal actions in settings where the optimal policy is multi-modal (mao2024diffusion, ), and CFG often fails to generalize when the conditioning signal diverges from those seen during training, resulting in unreliable plans.
To address these challenges, recent high-performing diffusion planners have adopted a simple yet effective alternative—Monte Carlo sample selection (MCSS) (hansen2023idql, ; lu2025makes, ). MCSS generates $N$ candidate trajectories and selects the one with the highest estimated return using a separately learned critic. While empirically successful, MCSS faces three primary limitations: (i) it incurs significant computational overhead at evaluation time due to the need to sample and score multiple trajectories, (ii) it does not guarantee the sampling of high-quality trajectories in the dataset, potentially overlooking valuable samples with high estimated value but low sampling likelihood, and (iii) increasing $N$ to improve coverage can lead to selecting out-of-distribution trajectories with spuriously high critic estimates.
Figure 1: Comparison between the behavior-cloned diffusion model and Prior Guidance. (Left) The behavior diffusion model samples latent noise $\mathbf{x}_{T}$ from a standard Gaussian and generates trajectories $\mathbf{x}_{0}$ via a learned denoising process $\tilde{\pi}_{\beta}(\mathbf{x}_{0}|\mathbf{x}_{T},\mathbf{s})$, approximating the dataset distribution. (Right) Prior Guidance instead samples $\mathbf{x}_{T}$ from a learned prior $p_{\psi}(\mathbf{x}_{T}|\mathbf{s})$ concentrated in high-value regions, leading to improved trajectory generation through the same denoising process.
Rather than guiding a pretrained diffusion model at inference time, an alternative approach is to directly train the diffusion model to generate high-value behaviors by maximizing a learned critic under behavior regularization (wang2022diffusion, ; ding2023consistency, ; ada2024diffusion, ; zhang2024entropy, ). However, this training objective typically incurs substantial computational overhead, as it requires backpropagation through the entire denoising process. Consequently, these efforts have primarily focused on developing diffusion policies rather than diffusion planners. Moreover, since diffusion models do not offer tractable density evaluation, most existing methods have relied on simple forms of behavior regularization, such as incorporating auxiliary behavior cloning losses.
To this end, we propose Prior Guidance (PG), a novel guided sampling method that trains the prior distribution that decodes into high-value trajectories. After the training of a diffusion model through behavior cloning, we replace the standard Gaussian prior with a learnable parametrized prior distribution. This learnable prior is optimized using a behavior-regularized objective, encouraging the sampling of high-value trajectories through the diffusion model while remaining close to the behavior distribution (See Figures 1 and 2).
PG effectively addresses the limitations of previous methods. In contrast to MCSS, PG generates a single trajectory during inference, significantly reducing computational overhead. Compared to end-to-end training of the diffusion model with a behavior-regularized objective, PG is more efficient, as it avoids backpropagation through the full denoising process and requires optimizing only a small set of parameters. Furthermore, by modeling the target prior as a simple Gaussian distribution, PG enables more expressive and analytically tractable behavior regularization, achieving a more favorable trade-off between maximizing potential performance and mitigating the risk of out-of-distribution action selection. Empirically, PG outperforms existing diffusion-based methods and achieves state-of-the-art performance on long-horizon tasks in the D4RL offline RL benchmark suite (fu2020d4rl, ).
Figure 2: Visualization of trajectory generation using three different guided sampling methods on maze2d-large-v1. Each method produces 20 planned trajectories. For MCSS, each trajectory is selected from a set of 10 candidates ($N=10$); the unselected candidates are visualized in gray.
2 Preliminaries
Diffusion Probabilistic Models
Diffusion models (ho2020denoising, ) are a class of generative models based on a Markovian forward–reverse process. The forward diffusion process gradually corrupts a data sample $\mathbf{x}_{0}$ by adding Gaussian noise over $T$ steps, defined as $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$, where $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$ and $\beta_{t}\in(0,1)$ is a predefined variance schedule. The reverse process is a parameterized Markov chain starting from a standard Gaussian prior $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$, defined as $p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$ and both $\mu_{\theta}$ and $\Sigma_{\theta}$ are predicted by a neural network.
The model is trained by minimizing a variational upper bound on the negative log-likelihood. In practice, a simplified objective is widely used, where the model learns to predict the added noise:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{0},\epsilon,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right\|^{2}\right],$ (1)
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. This training objective corresponds to denoising score matching across multiple noise levels. Once trained, the model can generate new samples by iteratively following the reverse process defined above. Alternatively, song2020denoising proposed another generative process that shares the same marginal distribution, defined by
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(\mathbf{x}_{t},t)+\sigma_{t}\bm{\epsilon},$
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{t}$ is a tunable parameter that controls the stochasticity of the generation. When $\sigma_{t}=0$, the generative process becomes entirely deterministic, resulting in the Denoising Diffusion Implicit Model (DDIM) sampling scheme.
Offline Reinforcement Learning
We assume the reinforcement learning problem under a Markov Decision Process (MDP). The MDP can be represented as a tuple $\mathcal{M}=\{S,\penalty 10000\ A,\penalty 10000\ P,\penalty 10000\ R,\penalty 10000\ \gamma,\penalty 10000\ p_{0}\}$, where $S$ is the state space, $A$ is the action space, $P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a}):S\times A\rightarrow\Delta(S)$ is a transition dynamics, $R(\mathbf{s},\mathbf{a}):S\times A\rightarrow\mathbb{R}$ is a reward function, $p_{0}\in\Delta(S)$ is an initial state distribution and $\gamma\in[0,1)$ is a discount factor. Offline RL aims to train a parameterized policy $\pi_{\psi}(\mathbf{a}|\mathbf{s})$, which maximizes the expected discounted return from a dataset $\mathcal{D}$ without environment interactions.
A common strategy in offline RL is to optimize policies via a behavior-regularized objective: $\max_{\psi}\mathbb{E}_{\mathbf{s}\sim\mathcal{D},\mathbf{a}\sim\pi_{\psi}(\cdot|\mathbf{s})}\left[Q(\mathbf{s},\mathbf{a})-\alpha f\left(\frac{\pi_{\psi}(\mathbf{a}|\mathbf{s})}{\pi_{\beta}(\mathbf{a}|\mathbf{s})}\right)\right]$, where $Q$ denotes the action-value function of policy $\pi_{\psi}$, $\pi_{\beta}$ is the behavior policy, and $f(\cdot)$ is a regularization function (e.g., for $f$-divergence) (wu2019behavior, ; xu2023offline, ; peng2019advantage, ; nair2020awac, ; chen2023score, ). However, when employing a diffusion policy or planner to model $\pi_{\psi}$, directly evaluating the density $\pi_{\psi}(\mathbf{a}|\mathbf{s})$ becomes intractable, making it challenging to apply sophisticated regularization functions. As a result, prior works have instead resorted to using a diffusion behavior cloning loss (1) as a surrogate for behavior regularization (wang2022diffusion, ; ding2023consistency, ; ada2024diffusion, ; zhang2024entropy, ).
Diffusion Planner
Diffusion-based planners (janner2022planning, ; ajay2022conditional, ; liang2023adaptdiffuser, ; li2023hierarchical, ; chen2024simple, ; lu2025makes, ) generate a trajectory of length $H$ starting from the current time step $\tau$. While various design choices are available, in this work, we largely follow the recommendations of lu2025makes , which are based on comprehensive experimental analysis. Specifically, a diffusion planner is trained to generate $H$-length sequences of current and future states $\mathbf{x}^{(\tau)}=[\mathbf{s}^{(\tau)},\mathbf{s}^{(\tau+m)},\dots,\mathbf{s}^{(\tau+(H-1)m)}]$ with the planning stride of $m$, where $\tau$ denotes the environment timestep. To bias the generation toward high-return behaviors, a trajectory critic $V(\mathbf{x}^{(\tau)})=\mathbb{E}\left[\sum_{h=0}^{\infty}\gamma^{h}r^{(\tau+h)}|\mathbf{x}^{(\tau)}\right]$ is estimated and utilized during planning. This critic can be either the behavior value function or the target value function. After producing a high-value state sequence $\mathbf{x}^{(\tau)}$, a separately trained inverse dynamics model is employed to recover the corresponding action $\mathbf{a}^{(\tau)}$ from the predicted sequence. For simplicity, we overload the notation and denote the diffusion planner-based policy as $\pi(\mathbf{x}|\mathbf{s})$.
A detailed discussion of related work is provided in the Appendix A.
3 Diffusion Planning with Prior Guidance
Figure 3: Predicted state sequences in maze2d-large-v1 based on MCSS varying the number of candidate trajectories $N$. When NN is small ($N=1$), MCSS often fails to sample trajectories with high value estimates, as the likelihood of capturing an optimal trajectory within a limited set is low. Conversely, when NN is excessively large ($N=500$ or 50005000), the method tends to select OOD trajectories with inflated value estimates.
Limitations of Monte Carlo Sample Selection (MCSS)
Monte Carlo Sample Selection (MCSS), which selects the best sample from $N$ candidates based on a separately trained value function, has been shown to achieve superior performance and has been widely adopted in recent diffusion-based offline RL (mao2024diffusion, ; hansen2023idql, ; lu2025makes, ). Nonetheless, as discussed above, MCSS also exhibits significant limitations: (i) it incurs substantial computational overhead during inference, as it requires sampling multiple candidates from diffusion models, which are already expensive to sample from, (ii) it does not guarantee coverage of high-value samples within the behavioral distribution, and (iii) while a sufficiently large $N$ is necessary for effective guidance, excessively large $N$ increases the risk of selecting OOD samples with overestimated value estimates, leading to performance degradation. [^1] A notable variant is SfBC (chen2022offline, ), which performs resampling by assigning weights proportional to $\exp(\alpha Q(\mathbf{s},\mathbf{a}))$ over a set of $N$ candidate actions generated by a diffusion policy. This weighting scheme implicitly induces KL-behavior regularization and mitigates limitation (iii). However, it introduces an additional hyperparameter $\alpha$, increasing the complexity of hyperparameter tuning. Moreover, due to other unresolved limitations, the method continues to incur significant inference-time computational overhead. Figure 3 illustrates these limitations.
3.1 Prior Guidance
To mitigate these shortcomings, we aim to adopt a behavior regularization framework in providing guidance for diffusion planners. Concretely, the objective is formalized as:
$\max_{\psi}\mathbb{E}_{\mathbf{s}\sim\mathcal{D},\mathbf{x}_{0}\sim\pi_{\psi}(\cdot|\mathbf{s})}\left[V(\mathbf{x}_{0})-\alpha f\left(\frac{\pi_{\psi}(\mathbf{x}_{0}|\mathbf{s})}{\pi_{\beta}(\mathbf{x}_{0}|\mathbf{s})}\right)\right],$ (2)
where $\mathbf{x}_{0}$ denotes a fully denoised trajectory generated by the diffusion model. This objective promotes the generation of high-value trajectories while constraining them to stay close to the dataset. However, directly optimizing this objective with diffusion models poses significant challenges: (i) it requires expensive backpropagation through the iterative denoising process, which becomes particularly costly for high-dimensional diffusion planners; and (ii) evaluating $\pi_{\psi}(\mathbf{x}_{0}|\mathbf{s})$ is non-trivial, as diffusion models do not admit tractable density estimation.
We now introduce Prior Guidance, a method for guiding diffusion models with a learnable prior distribution, which is free from the abovementioned challenges. Let $\tilde{\pi}_{\beta}$ denote a diffusion planner trained via behavior cloning. Given a state $\mathbf{s}$, the model generates a trajectory $\mathbf{x}_{0}$ by first sampling an initial noise $\mathbf{x}_{T}\sim p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$, followed by a denoising process. Assuming a DDIM sampling, the conditional density $\tilde{\pi}_{\beta}(\mathbf{x}_{0}|\mathbf{s})$ can be expressed as:
$\tilde{\pi}_{\beta}(\mathbf{x}_{0}|\mathbf{s})$ $=\mathbb{E}_{p(\mathbf{x}_{T})}\left[\delta(\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T}))\right],$
where $\delta(\cdot)$ denotes the Dirac delta function, and $g_{\mathbf{s}}$ is the deterministic denoising operator of the behavior diffusion model $\tilde{\pi}_{\beta}$ for state $\mathbf{s}$.
song2020denoising has demonstrated that interpolations in the latent space $\mathbf{x}_{T}$ yield smooth and semantically meaningful trajectories in the sample space, indicating that $\mathbf{x}_{T}$ captures the semantic structure of $\mathbf{x}_{0}$. This observation motivates shifting where the learning happens from the diffusion model to the prior distribution. Accordingly, we propose to learn a parameterized Gaussian prior $p_{\psi}(\mathbf{x}_{T}|\mathbf{s})$, while keeping the denoising operator $g_{\mathbf{s}}$ fixed. The trajectory distribution after planning $\pi_{\psi}$ can then be reformulated as:
$\pi_{\psi}(\mathbf{x}_{0}|\mathbf{s})=\mathbb{E}_{p_{\psi}(\mathbf{x}_{T}|\mathbf{s})}\left[\delta(\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T}))\right].$
When a sufficiently large number of discretization steps are employed, the denoising operator $g_{\mathbf{s}}$ can be viewed as an approximate, bijective ODE solver. This bijection allows us to re-express previously intractable behavior-regularized objective of the diffusion model in terms of a tractable objective over prior distributions:
Proposition 1.
Assume that DDIM sampling employs a sufficiently large number of discretization steps ensuring that the mapping $\mathbf{x_{0}}=g_{\mathbf{s}}(\mathbf{x}_{T})$ is bijective. If the target trajectory density $\pi_{\psi}$ is parametrized by placing the prior $p_{\psi}(\mathbf{x}_{T}|\mathbf{s})$ while keeping the denoising process $g_{\mathbf{s}}$ fixed, the behavior-regularized objective (2) is equivalent to:
$\max_{\psi}\mathbb{E}_{\mathbf{s}\sim\mathcal{D},\mathbf{x}_{T}\sim p_{\psi}(\cdot|\mathbf{s})}\left[V\left(g_{\mathbf{s}}\left(\mathbf{x}_{T}\right)\right)-\alpha f\left(\frac{p_{\psi}(\mathbf{x}_{T}|\mathbf{s})}{p(\mathbf{x}_{T})}\right)\right].$ (3)
The proof is in the Appendix B. Note that the behavior regularization term $f(\cdot)$ is now evaluated between two Gaussian distributions ($p_{\psi}$ and $p$), rather than between intractable diffusion model densities, which enables closed-form computation under several popular formulations of $f$.
Avoiding Backpropagation Through the Denoising Process
The reformulated objective (3) still depends on $g_{\mathbf{s}}$, requiring costly backpropagation through the denoising process. To avoid this, we introduce a value function $\bar{V}$ that operates directly in the latent prior space $\mathbf{x}_{T}$ rather than on the denoised output $\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T})$. Due to the bijectivity of $g_{\mathbf{s}}$ and the smoothness of the latent space induced by the diffusion model, the latent value function $\bar{V}$ is well-defined and straightforward to learn. It is trained by minimizing the following objective:
$\min_{\phi}\mathbb{E}_{\mathbf{x}_{T}\sim p_{\psi}(\cdot|\mathbf{s})}\left[\left(\bar{V}_{\phi}(\mathbf{x}_{T})-V\left(g_{\mathbf{s}}(\mathbf{x}_{T})\right)\right)^{2}\right],$ (4)
where $V(g_{\mathbf{s}}(\mathbf{x}_{T}))$ serves as the target signal. Once trained, the latent space value function $\bar{V}_{\phi}$ is used to replace the original critic in the behavior-regularized objective (3):
$\max_{\psi}\mathbb{E}_{\mathbf{s}\sim\mathcal{D},\mathbf{x}_{T}\sim p_{\psi}(\cdot|\mathbf{s})}\left[\bar{V}_{\phi}(\mathbf{x}_{T})-\alpha f\left(\frac{p_{\psi}(\mathbf{x}_{T}|\mathbf{s})}{p(\mathbf{x}_{T})}\right)\right].$ (5)
This training scheme completely eliminates the need for gradient flow through the denoising operator $g_{\mathbf{s}}$. Gradients are propagated only through the prior distribution $p_{\psi}$ and the value function $\bar{V}_{\phi}$, while $g_{\mathbf{s}}$ remains fixed. Consequently, PG circumvents the major computational bottleneck associated with directly optimizing diffusion models for high-value behavior generation, thereby enabling its application to diffusion planners, whereas previous approaches were constrained to diffusion policies due to their computational overhead. In practice, we alternate between minimizing the regression loss in Eq. (4) and optimizing the prior via Eq. (5). In Section 4.2, we qualitatively analyze the latent value function $\bar{V}$ and show that it accurately captures the intended value structure.
Figure 4: Training and inference time comparison. (Left) Training time until convergence for DQL (planner), PG with and without backpropagation through the denoising process across different environments. DQL (planner) denotes a baseline where the diffusion model is directly trained via Eq. (2); due to the intractability of computing the model’s density, Eq. (1) is used as a surrogate. (Right) Inference time for MCSS and PG. The inference time is measured on the antmaze-medium-play-v2, with MCSS evaluated under different numbers of sampled trajectories.
Figure 5: Average normalized score comparison between MCSS with varying numbers of samples NN and PG across four D4RL tasks. PG consistently matches or outperforms MCSS while requiring significantly fewer samples. Full results are provided in the Appendix D.
3.2 Does PG Address the Limitations of MCSS?
We empirically validate whether PG indeed resolves the limitations of MCSS, as claimed above. To ensure a fair comparison, all other experimental settings were held constant when comparing PG and MCSS.
(1) Reduced computational cost during inference. MCSS requires sampling multiple candidate trajectories at each environment step and selecting the one with the highest estimated value using a critic. Since sampling from a diffusion model is inherently computationally expensive, generating multiple candidates introduces substantial overhead, particularly in high-dimensional or long-horizon tasks. In contrast, PG generates a single trajectory starting from a learned prior, thereby avoiding this computational burden. Although PG introduces the additional cost of training the prior, it circumvents the need for costly backpropagation through the full denoising process by training a latent space value function, enabling efficient training.
As shown in Figure 4, PG w/o Bp—our proposed method that removes backpropagation through the denoising process—significantly reduces training time compared to PG w/ Bp. Moreover, PG achieves substantially lower inference time than MCSS, especially when the number of sampled trajectories $N$ is large, highlighting its superior computational efficiency. We also report the training time of Diffusion Q-learning when implemented as a planner-based method, denoted as DQL (planner). As expected, DQL (planner) is significantly more time-consuming to train than PG, taking up to ten times longer in the AntMaze domain. More experimental details for DQL (planner) are provided in the Appendix C.1.
(2) Tractable and stable behavior regularization. MCSS lacks an explicit mechanism to guide sampling toward trajectories aligned with the behavioral distribution, often resulting in OOD trajectory selections with spuriously high value estimates. PG addresses this limitation by performing regularization directly in the prior space, where the divergence between the learned prior $p_{\psi}$ and the standard Gaussian prior $p$ admits a closed-form expression. This enables stable and analytically tractable behavior regularization, which was previously infeasible in diffusion models due to the intractability of their output densities.
Empirical results presented in Figure 5 corroborate this advantage: while MCSS exhibits a trade-off between trajectory optimality and OOD avoidance, PG with appropriate behavior regularization consistently outperforms MCSS with the best performing $N$ across diverse domains.
Table 1: Normalized scores on the D4RL benchmark. For each dataset, the best-performing method is highlighted in red. Prior Guidance (PG) consistently achieves strong performance, particularly on long-horizon tasks, and outperforms DV* in nearly all environments. We compute the mean and standard error over 5 random seeds. Due to space constraints, the standard error for all algorithms are provided in the Appendix E.
Dataset Gaussian policies Diffusion policies Diffusion planners
BC CQL IQL SfBC DQL IDQL QGPO SRPO Diffuser AD $D$ HD DV* PG
Walker2d-M 6.6 79.2 78.3 77.9 87.0 82.5 86.0 84.4 79.6 84.4 82.5 84.0 79.5 82.3
Walker2d-M-R 11.8 26.7 73.9 65.1 95.5 85.1 84.4 84.6 70.6 84.7 75.0 84.1 83.5 83.7
Walker2d-M-E 6.4 111.0 109.6 109.8 110.1 112.7 110.7 114.0 106.9 108.2 108.8 107.1 109.0 109.4
Hopper-M 29.0 58.0 66.3 57.1 90.5 65.4 98.0 95.5 74.3 96.6 79.3 99.3 84.1 97.5
Hopper-M-R 11.3 48.6 94.7 86.2 101.3 92.1 96.9 101.2 93.6 92.2 100.0 94.7 91.3 91.3
Hopper-M-E 111.9 98.7 91.5 108.6 111.1 108.6 108.0 100.1 103.3 111.6 111.8 115.3 109.9 110.4
HalfCheetah-M 36.1 44.4 47.4 45.9 51.1 51.0 54.1 60.4 42.8 44.2 49.1 46.7 50.9 45.6
HalfCheetah-M-R 38.4 46.2 44.2 37.1 47.8 45.9 47.6 51.4 37.7 38.3 39.3 38.1 46.4 46.4
HalfCheetah-M-E 35.8 62.4 86.7 92.6 96.8 95.9 93.5 92.2 88.9 89.6 90.6 92.5 92.3 95.2
Average 31.9 63.9 77.0 75.6 87.9 82.1 86.6 87.1 77.5 83.3 81.8 84.6 83.0 84.6
Kitchen-M 47.5 51.0 51.0 45.4 62.6 66.5 52.5 51.8 65.0 71.7 73.3 74.6
Kitchen-P 33.8 49.8 46.3 47.9 60.5 66.7 55.7 55.5 57.0 73.3 76.2 88.0
Average 40.7 50.4 48.7 46.7 61.6 66.6 54.1 53.7 61.0 72.5 74.8 81.3
Antmaze-M-P 0.0 14.9 71.2 81.3 76.6 84.5 83.6 80.7 6.7 12.0 8.0 80.8 87.8
Antmaze-M-D 0.0 15.8 70.0 82.0 78.6 84.8 83.8 75.0 2.0 6.0 4.0 88.7 82.0 87.3
Antmaze-L-P 0.0 53.7 39.6 59.3 46.4 63.5 66.6 53.6 17.3 5.3 0.0 80.8 82.4
Antmaze-L-D 0.0 61.2 47.5 45.5 56.6 67.9 64.8 53.6 27.3 8.7 0.0 83.6 76.0 76.0
Average 0.0 36.4 57.1 67.0 64.6 75.2 74.7 65.7 13.3 8.0 3.0 79.9 83.4
Maze2D-U 3.8 5.7 47.4 73.9 57.9 113.9 135.1 128.4 133.8 139.2
Maze2D-M 30.3 5.0 34.9 73.8 89.5 121.5 129.9 135.6 144.1 159.5
Maze2D-L 5 12.5 58.6 74.4 90.1 123 167.9 155.8 197.6 195.2
Average 13.0 7.7 47.0 74.0 79.2 119.5 144.3 139.9 158.5 164.6
4 Experiments
Algorithm 1 Prior Guidance
Input: Dataset $D$, Hyperparameter $\alpha$
Require: Planner $g_{\mathbf{s}}$, Inverse Dynamics $\epsilon_{\omega}$, Critic $V$
Initialize: Prior $p_{\psi}$, Critic $\bar{V}_{\phi}$
1:// Training
2:for each iteration do
3:  Sample $\mathbf{s}\sim D$ and $\mathbf{x}_{T}\sim p_{\psi}(\cdot\mid\mathbf{s})$
4:  Update critic $\bar{V}_{\phi}$ using Eq. (4)
5:  Update prior $p_{\psi}$ using Eq. (5)
6:end for
7:// Execution
[…]
9:for environment step $\tau$ = 0, 1, … do
10:  Sample $\mathbf{x}_{T}\sim p_{\psi}(\cdot\mid\mathbf{s}^{(\tau)})$
11:  Generate $\mathbf{x}_{0}$ using $g_{\mathbf{s}}$
12:  Predict action $\mathbf{a}^{(\tau)}$ using $\epsilon_{\omega}(\mathbf{x}_{0})$
13:  $\mathbf{s}^{(\tau+1)}=\text{env.step}(\mathbf{a}^{(\tau)})$
14:end for
Practical Implementation
The target prior $p_{\psi}$ in PG is parameterized using a GRU (chung2014empirical, ) to capture the temporal dependencies in $\mathbf{x}$. Conditioned on the current state, the prior network outputs a sequence of mean and log standard deviation vectors that define independent Gaussian distributions over the latent noise dimensions. The current state is first provided as input to the GRU, and its output (prior to sampling) is fed back as input for subsequent steps. Detailed experimental settings are provided in the Appendix C.2.
Apart from the learnable prior, PG is built on top of Diffusion Veteran (DV) (lu2025makes, ), following their design choices. Due to the long training time of the original implementation, we re-implemented DV and refer to this version as DV*. PG inherits from DV* the planner $g_{\mathbf{s}}$, the inverse dynamics model $\epsilon_{\omega}$, and the critic $V$ as described in Algorithm 1; following DV, PG uses the behavior value function for its critic $V$. The pseudocode for DV* is provided in the Appendix D, along with a detailed comparison to the performance of the original DV implementation.
4.1 D4RL Benchmark
We conducted experiments on the D4RL offline RL benchmark (fu2020d4rl, ), which span a wide range of domains and dataset settings. We evaluate Prior Guidance (PG) across four representative tasks: MuJoCo, Kitchen, Antmaze, and Maze2D. In MuJoCo tasks, the dataset types are denoted as M (medium), R (replay), and E (expert); in Kitchen tasks, M (mixed) and P (partial); in Antmaze tasks, M (medium), L (large), P (play), and D (diverse); in Maze2D tasks, U (umaze), M (medium), and L (large). We benchmark against BC, CQL (kumar2020conservative, ), IQL (kostrikov2021offline, ), SfBC (chen2022offline, ), DQL (wang2022diffusion, ), IDQL (hansen2023idql, ), QGPO (lu2023contrastive, ), SRPO (chen2023score, ), Diffuser (janner2022planning, ), AD (liang2023adaptdiffuser, ), DD (ajay2022conditional, ), HD (chen2024simple, ), and DV (lu2025makes, ). The compared methods are categorized into Gaussian policies (CQL, IQL), diffusion policies (SfBC, DQL, IDQL, QGPO, SRPO), and diffusion planners (Diffuser, AD, DD, HD, DV). The normalized scores are summarized in Table 1.
As shown in Table 1, PG achieves state-of-the-art performance on tasks that require long-horizon decision making. By addressing the key limitations of MCSS, it significantly improves upon DV*. In MuJoCo tasks where complex decision making is not required, diffusion planners generally underperform compared to diffusion policies. Although PG also exhibits relatively lower performance in these tasks, it achieves the highest scores among diffusion planners. In long-horizon tasks such as Kitchen, AntMaze, and Maze2D, PG substantially outperforms all baselines across most tasks.
4.2 Analysis
Figure 6: Visualizations of trajectories $(\mathbf{x}_{0})$ and priors $(\mathbf{x}_{T})$ from the behavior diffusion model and Prior Guidance in maze2d-large-v1 (top) and antmaze-medium-diverse-v2 (bottom). We applied PCA (jolliffe2016principal, ) to project the high-dimensional trajectories into two dimensions, with the color of each sample representing its estimated value. We plotted 5,000 samples from the standard normal prior of the behavior diffusion model and 1,000 samples from the target prior of PG.
(a) Correlation between $\bar{V}_{\phi}(\mathbf{x}_{T})$ and $V(\mathbf{x}_{0})$ in walker2d and hopper environments.
Dataset PG w/ Bp PG w/o Bp
[…]
Total 415.0 413.9
(b) Average scores with and without backpropagation through the denoising process.
Figure 7: Experiments to evaluate whether the latent value function $\bar{V}$ is accurately trained.
Visualization on Latent Trajectories
We visualize the latent noise and corresponding trajectories from the behavior diffusion model and PG to verify whether the learned prior $p_{\psi}$ in PG accurately captures the noise $\mathbf{x}_{T}$ that is denoised into high-value trajectories $\mathbf{x}_{0}$. As shown in Figure 6, the prior of the behavior diffusion model spans a wide range of values, whereas the learned prior of PG predominantly samples noise that leads to high-value trajectories. This demonstrates that the latent space of the learned behavior diffusion model is structured sufficiently well to cluster high-value trajectories under a simple Gaussian prior. Although the target prior in PG remains close to the standard normal distribution due to behavior regularization, it still captures mostly high-value trajectories.
Consequence of Using Latent‐Space Value Function
To avoid backpropagation through the denoising process, PG trains a latent-space value function $\bar{V}$ directly on the prior noise $\mathbf{x}_{T}$ (4). Figure 7 presents our empirical evaluation of this design choice. In particular, Figure 7(a) depicts the correlation between learned latent values $\bar{V}_{\phi}(\mathbf{x}_{T})$ and ground-truth values $V(\mathbf{x}_{0})$ for the walker2d-medium-v2 and hopper-medium-v2 tasks. While a strong positive correlation is observed, a perfect learning of the mapping would yield a correlation of exactly $1$. The observed deviation reflects approximation errors, which could potentially affect performance when relying on the latent value function. However, our experiments show that these approximation errors do not lead to actual performance degradation. As shown in Figure 7(b), PG achieves nearly identical normalized scores on the D4RL benchmark with and without backpropagation through the denoising process, confirming that training $\bar{V}$ in latent space is sufficient. Full results corresponding to Figure 7(b) are provided in the Appendix G.
More Expressive Target Prior
Table 2: Average normalized scores of uni-modal and multi-modal Gaussian prior on MuJoCo and Antmaze tasks.
Dataset uni-modal multi-modal
[…]
To investigate the flexibility and representational capacity of PG, we extend our approach to use a multi-modal Gaussian mixture prior in place of the standard uni-modal Gaussian prior. This modification allows us to evaluate whether a more expressive prior distribution can yield performance gains. As shown in Table 2, the results are domain-dependent: in MuJoCo tasks, the uni-modal Gaussian prior slightly outperforms the multi-modal variant, whereas in AntMaze tasks, the multi-modal prior provides marginal improvements. These findings suggest that the benefit of a more expressive prior is more pronounced in complex, long-horizon environments such as AntMaze, while in simpler control tasks like MuJoCo, the latent space found by a diffusion model is sufficiently well structured to capture high-value trajectories only with a uni-modal gaussian distribution. Full results corresponding to Table 2 are provided in the Appendix G.
4.3 Ablation Studies
Regularization Function $f$
By deriving a closed‐form solution for behavior regularization, PG accommodates a variety of regularization functions $f$. To assess the impact of different choices of $f$, we performed an ablation study on the D4RL benchmark. We compared $f(x)=-\frac{\log x}{x}$ (KL-divergence), $f(x)=\log x$ (reverse KL-divergence), and $f(x)=\frac{(x-1)^{2}}{x}$ (Pearson $\chi^{2}$-divergence). In practice, we use KL-divergence in PG. Ablation results are provided in the Appendix H.
Behavior Regularization Coefficient $\alpha$
We present ablation studies evaluating the performance of PG across different values of the behavior regularization coefficient $\alpha$. In practice, we sweep $\alpha\in\{50.0,10.0,1.0,0.1,0.01,0.001\}$ in PG. Ablation results are provided in the Appendix H.
5 Conclusion
We introduced Prior Guidance, a guided sampling method for diffusion planner-based offline RL that replaces the standard Gaussian prior with a learnable, behavior-regularized distribution. Unlike previous guided sampling methods, PG avoids inference-time overhead, reduces the risk of out-of-distribution actions, and enables stable behavior regularization in closed form. Empirically, it achieves strong performance across diverse tasks, especially in long-horizon domains.
Limitations
Despite its effectiveness, PG has several limitations. First, it introduces additional complexity through the training of both a latent value function and a learnable prior. Second, the architectural choices for the prior network have not been thoroughly investigated and warrant deeper exploration. Third, our evaluation is limited to offline settings and does not consider high-dimensional observations like images. Finally, the bijective mapping between latent noise and trajectory assumed by DDIM does not hold in practice due to limited discretization steps; this could be addressed by using flow-matching (lipman2022flow, ; park2025flow, ), which we leave for future work.
6 Acknowledgements
This work was supported by the IITP(Institute of Information & Coummunications Technology Planning & Evaluation)-ITRC(Information Technology Research Center)(IITP-2025-RS-2024-00436857) grant funded by the Korea government(Ministry of Science and ICT) and by the IITP under the Artificial Intelligence Star Fellowship support program to nurture the best talents (IITP-2025-RS-2025-02304828) grant funded by the Korea government(MSIT).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Section 1
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in the Appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 4, Appendix C, Appendix D, Appendix F, Appendix G, Appendix H
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Abstract, Section 4, Appendix C
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Appendix F
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in the Appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Appendix C.3, Appendix E
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix C.2
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper does not have societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not release new datasets or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Section 4, Appendix D
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: Core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Related Works
Learning Priors in RL
Several prior works in reinforcement learning have utilized trainable prior distributions to achieve great success in multi-task learning [37] and skill learning [38, 39, 40, 41, 42]. [37] introduce a behavior prior framework that learns a distribution over trajectories to encode reusable movement and interaction patterns across tasks. [38] propose SPiRL, which leverages a learned skill prior to guide exploration and improve downstream task performance. [41] propose OPAL, which learns a latent prior over temporally extended primitive behaviors from offline data to improve downstream task performance in offline RL. [40] propose FIST, which learns a future-conditioned skill prior from offline data and adapts it for few-shot imitation of unseen long-horizon tasks. [39] propose SkiMo, which jointly learns a skill prior and a skill dynamics model from offline data to enable accurate long-horizon planning and improve downstream task performance. [42] propose NBDI, which learns a termination prior from a state-action novelty module to guide the execution of variable-length skills in downstream tasks.
While these methods learn priors over abstract skills or trajectories, they do not operate in the context of diffusion models. To the best of our knowledge, Prior Guidance (PG) is the first to replace the standard normal prior distribution of a diffusion model with a learnable prior in the context of reinforcement learning.
Learning Priors in Diffusion Models
Recent studies have explored methods to improve diffusion models by replacing the standard Gaussian prior with a more informative distribution. Notably, PriorGrad [43] proposes a data-dependent adaptive prior to accelerate training and improve sample quality in conditional generative tasks such as speech synthesis. While PriorGrad shares the motivation of improving the expressiveness and utility of the diffusion model’s prior distribution, PriorGrad applies prior learning to the forward process in conditional generative models and leverages exact statistics from input conditions to define an instance-specific non-learned prior. In contrast, PG learns a parameterized prior distribution over the latent space of a behavior-cloned diffusion model, specifically targeting the reverse denoising process for high-return trajectory generation in offline RL.
Behavior-Regularized Objectives in Offline RL
To mitigate policy degradation caused by overestimated value functions, a wide range of offline RL methods incorporate behavior-regularized objectives. [2] introduce BRAC, a unified framework that employs behavior regularization to stabilize policy learning. [26] propose Advantage-Weighted Regression (AWR), which formulates offline RL as supervised regression weighted by advantage estimates. [27] present AWAC, an actor-critic method that performs advantage-weighted supervised learning, enabling efficient offline pretraining and robust online fine-tuning. [44] propose OptiDICE, which optimizes a behavior-regularized objective over stationary distributions, avoiding the need for value bootstrapping or policy gradients. [4] introduce SQL and EQL, which leverage in-sample value regularization to address out-of-distribution (OOD) actions. [28] present SRPO, which directly regularizes the policy gradient using the score function of a pretrained diffusion model, thereby enabling behavior-regularized learning without sampling. [45] propose PORelDICE, which relaxes the positivity constraint in $f$-divergence regularization to better handle low-advantage samples. However, integrating behavior-regularized objectives into diffusion planners poses several challenges, such as requiring backpropagation through the iterative denoising process and making it difficult to handle a wide range of regularization functions in closed form.
Diffusion Models in Offline RL
Inspired by the success of diffusion models across a variety of domains [5, 6, 7, 8, 9, 10, 46], there has been growing interest in adapting diffusion models to offline RL. One line of research replaces conventional uni-modal Gaussian policies with expressive diffusion-based policies. For example, [20] propose Diffusion-QL, which combines conditional diffusion modeling with Q-learning guidance to improve policy expressiveness and regularization. [18] reinterpret IQL as an actor-critic algorithm and propose IDQL, which couples a generalized IQL critic with a diffusion-based behavior model. [33] introduce QGPO, which enables exact energy-guided diffusion sampling by aligning sampled trajectories with the energy landscape defined by a Q-function.
Another line of work leverages the temporal modeling capabilities of diffusion models to develop trajectory-level planners for long-horizon tasks. [11] introduce Diffuser, which formulates planning as conditional trajectory generation and enables flexible behavior synthesis via classifier-guided sampling. [12] propose Decision Diffuser, which formulates decision-making as return-conditioned generative modeling without relying on value estimation. [13] present AdaptDiffuser, a self-evolving planner that iteratively improves its performance through reward-guided trajectory generation and model fine-tuning. [14] propose HDMI, a hierarchical diffusion model that generates subgoals conditioned on returns and performs trajectory denoising toward each subgoal. [15] propose Hierarchical Diffuser, a two-level architecture that combines jumpy subgoal generation with fine-grained trajectory refinement. Finally, [19] conduct a comprehensive empirical study of diffusion planners and introduce Diffusion Veteran, a simple baseline derived from key design principles such as sampling strategy, backbone architecture, and action generation scheme. However, existing guided sampling methods in diffusion planners suffer from several limitations—such as suboptimal sample selection, distributional shift, and high computational cost—highlighting the need for novel guidance approaches that can more effectively balance performance and efficiency.
Concurrent Works
A concurrent work, DSRL [47], introduces a method that steers a behavior-cloned diffusion policy by optimizing a learnable prior through reinforcement learning. While this idea of guiding generative models via a learned prior shares the same high-level goal as our approach, there exist both conceptual similarities and key methodological differences between DSRL and our PG framework.
Both DSRL and PG enhance decision quality by learning a prior distribution over the noise space $(\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}))$ that serves as input to the diffusion model’s denoising process. In both methods, the prior is trained to maximize a value function defined over $\mathbf{x}_{T}$, which estimates the value of the corresponding denoised sample $\mathbf{x}_{0}$. This shared mechanism allows both methods to bias sampling toward high-value regions of the data or trajectory space.
However, there are fundamental differences. DSRL adopts a diffusion policy that directly maps noise to individual actions, whereas PG employs the diffusion planner that produces entire state trajectories and subsequently reconstructs executable actions via an inverse dynamics model. This trajectory-level formulation makes PG particularly effective in long-horizon or sparse-reward environments, where modeling temporal dependencies is critical for performance. Moreover, PG provides a theoretical foundation absent in DSRL. Specifically, we prove in Proposition 1 that, under sufficiently fine discretization, the optimization objective for the diffusion model can be equivalently reformulated as an optimization over the latent prior distribution. This equivalence arises from the bijective nature of the DDIM sampling process, which allows for an exact mapping between the data space and the latent prior. Finally, PG explicitly addresses the challenge of out-of-distribution (OOD) samples in offline reinforcement learning. While DSRL assumes the diffusion model generates only in-distribution actions and omits behavior constraints, our experiments (Figures 3 and 5) reveal that diffusion models trained on offline datasets can still produce OOD samples that harm performance. To mitigate this issue, PG introduces behavior regularization by imposing an $f$-divergence penalty between the learned prior and the standard normal distribution. This explicitly penalizes deviation from the behavior distribution and induces regularization at the prior level.
Appendix B Proof of Proposition 1
See 1
Proof.
Since $\mathbf{x}_{0}=g_{s}(\mathbf{x}_{T})$ is bijective mapping between $\mathbf{x}_{T}$ and $\mathbf{x}_{0}$, we can apply the change-of-variables formula:
$d\mathbf{x}_{0}=d\mathbf{x}_{T}\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|.$
Using this, we compute the conditional density $\tilde{\pi}_{\beta}(\mathbf{x}_{0}|\mathbf{s})$ as follows:
$\tilde{\pi}_{\beta}(\mathbf{x}_{0}|\mathbf{s})$ $=\int_{\mathbf{x}_{T}}p(\mathbf{x}_{T})\cdot\delta(\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T}))d\mathbf{x}_{T}$
$=\int_{\mathbf{x}_{0}}p(\mathbf{x}_{T})\cdot\delta(\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T}))d\mathbf{x}_{0}\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|^{-1}$
$=p(\mathbf{x}_{T})\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|^{-1}.$
Similarly, the density $\pi_{\psi}(\mathbf{x}_{0}|\mathbf{s})$ can be obtained in the same way:
$\pi_{\psi}(\mathbf{x}_{0}|\mathbf{s})$ $=\int_{\mathbf{x}_{T}}p_{\psi}(\mathbf{x}_{T}|\mathbf{s})\cdot\delta(\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T}))d\mathbf{x}_{T}$
$=\int_{\mathbf{x}_{0}}p_{\psi}(\mathbf{x}_{T}|\mathbf{s})\cdot\delta(\mathbf{x}_{0}=g_{\mathbf{s}}(\mathbf{x}_{T}))d\mathbf{x}_{0}\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|^{-1}$
$=p_{\psi}(\mathbf{x}_{T}|\mathbf{s})\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|^{-1}.$
Accordingly, we can reformulate the objective 2 as follows:
(2) $=\max_{\psi}\mathbb{E}_{\mathbf{s}\sim D,\mathbf{x}_{0}\sim\pi_{\psi}(\cdot|\mathbf{s})}\left[V(\mathbf{x}_{0})-\alpha f\left(\frac{\pi_{\psi}(\mathbf{x}_{0}|\mathbf{s})}{\tilde{\pi}_{\beta}(\mathbf{x}_{0}|\mathbf{s})}\right)\right]$
$=\max_{\psi}\mathbb{E}_{\mathbf{s}\sim D,\mathbf{x}_{T}\sim p_{\psi}(\cdot|\mathbf{s})}\left[V\left(g_{\mathbf{s}}(\mathbf{x}_{T})\right)-\alpha f\left(\frac{p_{\psi}(\mathbf{x}_{T}|\mathbf{s})\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|^{-1}}{p(\mathbf{x}_{T})\cdot\left|\det\left(\frac{\partial g_{\mathbf{s}}(\mathbf{x}_{T})}{\partial\mathbf{x}_{T}}\right)\right|^{-1}}\right)\right]$
$=\max_{\psi}\mathbb{E}_{\mathbf{s}\sim D,\mathbf{x}_{T}\sim p_{\psi}(\cdot|\mathbf{s})}\left[V\left(g_{\mathbf{s}}(\mathbf{x}_{T})\right)-\alpha f\left(\frac{p_{\psi}(\mathbf{x}_{T}|\mathbf{s})}{p(\mathbf{x}_{T})}\right)\right]$
Appendix C Experimental Details
C.1 Experiment Details of Figure 4
Algorithm 2 DQL (planner)
Input: Dataset $D$, Hyperparameter $\alpha$
Require: Inverse Dynamics $\epsilon_{\omega}$, Critic $V$
Initialize: Planner $\pi_{\psi}$
1:// Training
2:for each iteration do
3:  Sample $\mathbf{s}\sim D$ and $\mathbf{x}_{T}\sim p(\mathbf{x}_{T})$
4:  Update planner $\pi_{\psi}$ using Eq. (6)
5:end for
6:// Execution
[…]
8:for environment step $\tau$ = 0, 1, … do
9:  Sample $\mathbf{x}_{0}\sim\pi_{\psi}(\cdot|\mathbf{s})$
10:  Predict action $\mathbf{a}^{(\tau)}$ using $\epsilon_{\omega}(\mathbf{x}_{0})$
11:  $\mathbf{s}^{(\tau+1)}=\text{env.step}(\mathbf{a}^{(\tau)})$
12:end for
In Figure 4, we introduce a new baseline referred to as DQL (planner), which directly trains the diffusion model using a behavior-regularized objective (2). Specifically, it employs the diffusion behavior cloning loss (1) as a surrogate for behavior regularization, resulting in the following objective:
$\max_{\psi}\mathbb{E}_{\mathbf{s}\sim D,\mathbf{x}_{0}\sim\pi_{\psi}(\cdot|\mathbf{s})}\left[V(\mathbf{x}_{0})\right]-\alpha\mathcal{L}(\psi),$ (6)
where $\mathcal{L}(\cdot)$ denotes the behavior cloning loss defined in Eq.(1). This approach can be seen as the planner version of Diffusion Q-Learning (DQL)[20], and we refer to it as DQL (planner) throughout the paper. The corresponding pseudo-code is provided in Algorithm 2. We measure the training time of DQL (planner) after aligning all other settings with those of Prior Guidance to ensure a fair comparison.
In Figure 4, we compare the training time of PG with and without backpropagation through the denoising process to specifically assess the efficiency of learning the latent value function. To this end, we exclude the time required to train the planner $g_{\mathbf{s}}$, inverse dynamics model $\epsilon_{\omega}$, and critic $V$, which are shared across both variants in Algorithm 1. Likewise, for DQL (planner), we report only the time taken to optimize the diffusion model for value maximization, excluding the time for training the inverse dynamics, critic, and the behavior cloning loss—which aligns with the planner training time in PG. This setup enables us to assess how much training time can be saved through latent value function learning.
C.2 Experimental Settings
We evaluate Prior Guidance on the D4RL offline RL benchmark. For MuJoCo tasks, we report the average performance over 10 evaluation trajectories for each of 5 independently trained models. For Kitchen, AntMaze, and Maze2D tasks, we average over 100 evaluation trajectories for each of 5 independently trained models. In Prior Guidance, we adopt the same experimental settings for the planner $g_{\mathbf{s}}$, inverse dynamics model $\epsilon_{\omega}$, and critic $V$ as used in Diffusion Veteran (DV) [19]. We conducted all experiments using four NVIDIA RTX 4090 GPUs.
C.3 Statistical Details
For all experiments, we report the mean and standard error of normalized scores computed over 5 training seeds. Additionally, in Figure 5, Figure 7(b), and Table 2, we report the mean and standard error $\bar{\sigma}_{\text{avg}}$ of average normalized scores across multiple environments. The standard error $\bar{\sigma}_{\text{avg}}$ is computed using the error propagation formula based on the per-environment standard errors $\bar{\sigma}_{i}$, each estimated from 5 training seeds. Specifically,
$\bar{\sigma}_{\text{avg}}=\sqrt{\sum_{i=1}^{n}\bar{\sigma}_{i}^{2}}\bigg/n,\quad\text{where}\penalty 10000\ \bar{\sigma}_{i}=\frac{\sigma_{i}}{\sqrt{k}},$
with $k=5$ denoting the number of training seeds, $n$ the number of environments, and $\sigma_{i}$ the standard deviation for environment $i$.
Appendix D Diffusion Veteran (DV)
The pseudo-code for Diffusion Veteran (DV), which is identical to that of DV, is provided in Algorithm 3.
Algorithm 3 Diffusion Veteran (DV)
Input: Dataset $D$, Planning horizon $H$, Planning stride $m$, Candidate num NN
Initialize: Planner $g_{\mathbf{s}}$, Inverse Dynamics $\epsilon_{\omega}$, Critic $V$
1:Compute discounted returns: $R^{(\tau)}=\sum_{h=0}^{\infty}\gamma^{h}r^{(\tau+h)}\penalty 10000\ \penalty 10000\ \text{for every step }\tau$
2:// Training
3:for each iteration do
4:  Sample $\mathbf{s}^{(\tau)},\mathbf{s}^{(\tau+m)},\dots,s^{(\tau+(H-1)m))}$ and $\mathbf{a}^{(\tau)},R^{(\tau)}$ from $D$
5:  Update planner $g_{\mathbf{s}}$ using Eq. (1) with $\mathbf{s}^{(\tau)}$ as input, $\mathbf{s}^{(\tau)},\dots,\mathbf{s}^{(\tau+(H-1)m))}$ as target output
6:  Update inverse dynamics $\epsilon_{\omega}$ using Eq. (1) with $\mathbf{s}^{(\tau)},\mathbf{s}^{(\tau+m)}$ as input, $\mathbf{a}^{(\tau)}$ as target output
7:  Update critic $V$ using MSE Loss with $\mathbf{s}^{(\tau)},\dots,\mathbf{s}^{(\tau+(H-1)m))}$ as input, $R^{(\tau)}$ as target output
8:end for
9:// Execution (MCSS)
[…]
11:for environment step $\tau$ = 0, 1, … do
12:  Randomly generate NN candidate trajectories using $g_{\mathbf{s}}$ while fixing the first state as $\mathbf{s}^{(\tau)}$
13:  Select the trajectory with the highest estimated value $V$ among the NN candidates
14:  Generate $\mathbf{a}^{(\tau)}$ from the best trajectory using the inverse dynamics model $\epsilon_{\omega}$
15:  $\mathbf{s}^{(\tau+1)}=\text{env.step}(\mathbf{a}^{(\tau)})$
16:end for
Full results corresponding to Figure 5, along with the comparison between the original implementation of DV and Prior Guidance, are provided in Table 3.
Table 3: Normalized scores on the D4RL benchmark. we compare Prior Guidance (PG) with the original implementation DV and DV* under different NNs. We compute the mean and the standard error over 5 random seeds.
DV* DV PG
NN 1 5 10 50 (default) 500 50 1
walker2d-medium-v2 70.8 $\pm$ 2.2 81.1 $\pm$ 0.2 80.6 $\pm$ 0.3 79.5 $\pm$ 0.3 79.5 $\pm$ 0.7 82.8 $\pm$ 0.5 82.3 $\pm$ 0.2
walker2d-medium-replay-v2 40.3 $\pm$ 1.8 82.5 $\pm$ 0.9 84.1 $\pm$ 0.4 83.5 $\pm$ 2.5 85.9 $\pm$ 0.4 85.0 $\pm$ 0.1 83.7 $\pm$ 1.0
walker2d-medium-expert-v2 96.1 $\pm$ 2.1 108.9 $\pm$ 0.1 109.0 $\pm$ 0.1 109.0 $\pm$ 0.1 108.9 $\pm$ 0.2 109.2 $\pm$ 0.0 109.4 $\pm$ 0.1
hopper-medium-v2 47.2 $\pm$ 0.9 67.4 $\pm$ 2.8 70.1 $\pm$ 3.5 84.1 $\pm$ 2.5 91.5 $\pm$ 1.7 83.6 $\pm$ 1.2 97.5 $\pm$ 0.6
hopper-medium-replay-v2 52.4 $\pm$ 2.1 91.2 $\pm$ 0.2 91.0 $\pm$ 0.1 91.3 $\pm$ 0.1 91.3 $\pm$ 0.2 91.9 $\pm$ 0.0 91.3 $\pm$ 0.3
hopper-medium-expert-v2 76.2 $\pm$ 5.9 110.0 $\pm$ 0.3 110.2 $\pm$ 0.5 109.9 $\pm$ 1.0 110.6 $\pm$ 0.9 110.0 $\pm$ 0.5 110.4 $\pm$ 0.1
halfcheetah-medium-v2 41.4 $\pm$ 0.6 47.5 $\pm$ 0.1 49.2 $\pm$ 0.0 50.9 $\pm$ 0.1 52.3 $\pm$ 0.1 50.4 $\pm$ 0.0 45.6 $\pm$ 0.5
halfcheetah-medium-replay-v2 35.4 $\pm$ 1.4 43.2 $\pm$ 0.6 45.4 $\pm$ 0.2 46.4 $\pm$ 0.2 46.5 $\pm$ 0.6 45.8 $\pm$ 0.1 46.4 $\pm$ 0.4
halfcheetah-medium-expert-v2 82.3 $\pm$ 1.4 91.8 $\pm$ 0.2 91.6 $\pm$ 1.2 92.3 $\pm$ 0.6 87.9 $\pm$ 2.0 92.7 $\pm$ 0.3 95.2 $\pm$ 0.1
Average 60.2 80.4 81.2 83.0 83.8 83.5 84.6
kitchen-mixed-v0 52.4 $\pm$ 0.7 65.1 $\pm$ 1.2 70.3 $\pm$ 1.7 73.3 $\pm$ 0.6 67.5 $\pm$ 3.0 73.6 $\pm$ 0.1 74.6 $\pm$ 0.4
kitchen-partial-v0 53.4 $\pm$ 1.0 77.7 $\pm$ 2.0 80.6 $\pm$ 1.8 76.2 $\pm$ 4.3 66.8 $\pm$ 5.4 94.0 $\pm$ 0.3 88.0 $\pm$ 3.1
Average 52.9 71.4 75.5 74.8 67.2 83.8 81.3
antmaze-medium-play-v2 66.6 $\pm$ 3.8 86.4 $\pm$ 1.4 86.0 $\pm$ 1.2 80.8 $\pm$ 1.3 62.2 $\pm$ 3.2 89.0 $\pm$ 1.6 87.8 $\pm$ 1.5
antmaze-medium-diverse-v2 63.6 $\pm$ 2.0 72.4 $\pm$ 2.4 71.8 $\pm$ 4.0 82.0 $\pm$ 3.6 84.6 $\pm$ 1.6 87.4 $\pm$ 1.6 87.3 $\pm$ 1.7
antmaze-large-play-v2 74.8 $\pm$ 1.3 79.2 $\pm$ 2.0 80.6 $\pm$ 3.4 80.8 $\pm$ 2.8 81.0 $\pm$ 2.0 76.4 $\pm$ 2.0 82.4 $\pm$ 2.1
antmaze-large-diverse-v2 78.8 $\pm$ 2.6 78.2 $\pm$ 1.7 77.6 $\pm$ 3.0 76.0 $\pm$ 2.3 72.0 $\pm$ 1.4 80.0 $\pm$ 1.8 76.0 $\pm$ 1.9
Average 71.0 79.1 79.0 79.9 75.0 83.2 83.4
maze2d-umaze-v1 41.5 $\pm$ 3.4 97.1 $\pm$ 3.5 111.0 $\pm$ 2.7 133.8 $\pm$ 1.9 139.7 $\pm$ 1.5 136.6 $\pm$ 1.3 139.2 $\pm$ 1.3
maze2d-medium-v1 23.0 $\pm$ 3.5 95.3 $\pm$ 7.8 123.8 $\pm$ 5.9 144.1 $\pm$ 2.1 151.2 $\pm$ 2.6 150.7 $\pm$ 1.0 159.5 $\pm$ 0.8
maze2d-large-v1 54.6 $\pm$ 5.1 183.9 $\pm$ 5.3 190.8 $\pm$ 5.0 197.6 $\pm$ 4.2 201.8 $\pm$ 5.5 203.6 $\pm$ 1.4 195.2 $\pm$ 2.3
Average 39.7 125.4 141.9 158.5 164.2 163.6 164.6
Appendix E Full results of Table 1
Table 4: Normalized scores on the D4RL benchmark. We compute the mean and standard error over 5 random seeds.
Dataset Gaussian policies Diffusion policies Diffusion planners
BC CQL IQL SfBC DQL IDQL QGPO SRPO Diffuser AD $D$ HD DV* PG
[…]
Average 31.9 63.9 77.0 75.6 87.9 82.1 86.6 87.1 77.5 83.3 81.8 84.6 83.0 84.6
[…]
Average 40.7 50.4 48.7 46.7 61.6 66.6 54.1 53.7 61.0 72.5 74.8 81.3
[…]
Average 0.0 36.4 57.1 67.0 64.6 75.2 74.7 65.7 13.3 8.0 3.0 79.9 83.4
[…]
Average 13.0 7.7 47.0 74.0 79.2 119.5 144.3 139.9 158.5 164.6
Appendix F Hyperparameters
This section outlines the hyperparameters used to train Prior Guidance. The planner $g_{s}$, inverse dynamics model $\epsilon_{\omega}$, and critic $V$ all follow the same settings as those used in Diffusion Veteran. Detailed hyperparameter settings are provided in Table 5.
Table 5: Hyperparameter settings used for Prior Guidance on the D4RL benchmark.
Settings MuJoCo Kitchen Antmaze Maze2d
Batch Size 128 128 128 128
State-Action Generation Separate Separate Separate Separate
Advantage Weighting True False False False
Inverse Dynamic Diffusion Diffusion Diffusion Diffusion
Time Credit Assignment 0.997 0.997 1.0 1.0
Planner Net. Backbone Transformer Transformer Transformer Transformer [48]
Transformer Hidden 256 256 256 256
Transformer Block 2 2 2 8
Planner Solver DDIM DDIM DDIM DDIM [25]
Planner Sampling Steps 20 20 20 20
Planner Training Steps 1M 1M 1M 1M
Planner Temperature 1 1 1 1
Planner Learning Rate 2e-4 2e-4 2e-4 2e-4
Planner Optimizer AdamW AdamW AdamW AdamW [49]
Planning Horizon 4 32 32 40
Planning Stride 1 1 15 25
Inverse Dynamics Net. Backbone MLP MLP MLP MLP
Inverse Dynamics Hidden 256 256 256 256
Inverse Dynamics Solver DDPM DDPM DDPM DDPM [5]
Inverse Dynamics Sampling Steps 10 10 10 10
Inverse Dynamics Training Steps 1M 1M 1M 1M
Policy Temperature 0.5 0.5 0.5 0.5
Policy Learning Rate 3e-4 3e-4 3e-4 3e-4
Policy Optimizer AdamW AdamW AdamW AdamW
Value Learning Rate 3e-4 3e-4 3e-4 3e-4
Value Net. Backbone Transformer Transformer Transformer Transformer
Value Optimizer Adam Adam Adam Adam [50]
Prior Network GRU GRU GRU GRU [30]
Prior Learning Rate 3e-4 3e-4 3e-4 3e-4
Prior Optimizer AdamW AdamW AdamW AdamW
Prior Hidden 256 256 256 256
Latent Value Learning Rate 3e-4 3e-4 3e-4 3e-4
Latent Value Net. Backbone Transformer Transformer Transformer Transformer
Latent Value Optimizer Adam Adam Adam Adam
Behavior Regularization KL KL KL KL
In Table 15, we present ablation studies on the choice of behavior regularization function $f$. The corresponding hyperparameter $\alpha$ are listed in Table 6 and 7.
Table 6: $\alpha$ used for KL and Reverse KL-divergence on the D4RL benchmark.
$\alpha$ KL Reverse KL
walker2d-medium-v2 0.1 0.1
walker2d-medium-replay-v2 0.01 0.01
walker2d-medium-expert-v2 0.01 0.01
hopper-medium-v2 0.01 0.01
hopper-medium-replay-v2 0.01 0.001
hopper-medium-expert-v2 0.01 0.01
halfcheetah-medium-v2 0.001 0.0001
halfcheetah-medium-replay-v2 0.001 0.001
halfcheetah-medium-expert-v2 0.01 0.01
kitchen-mixed-v0 1.0 10.0
kitchen-partial-v0 1.0 10.0
antmaze-medium-play-v2 1.0 1.0
antmaze-medium-diverse-v2 1.0 0.1
antmaze-large-play-v2 50.0 1.0
antmaze-large-diverse-v2 10.0 10.0
maze2d-umaze-v1 1.0 10.0
maze2d-medium-v1 0.1 1.0
maze2d-large-v1 0.1 1.0
Table 7: $\alpha$ used for Pearson $\chi^{2}$-divergence on the D4RL benchmark.
$\alpha$ Pearson $\chi^{2}$
walker2d-medium-v2 1e-4
walker2d-medium-replay-v2 1e-7
walker2d-medium-expert-v2 1e-6
hopper-medium-v2 1e-8
hopper-medium-replay-v2 1e-6
hopper-medium-expert-v2 1e-5
halfcheetah-medium-v2 1e-4
halfcheetah-medium-replay-v2 1e-6
halfcheetah-medium-expert-v2 1e-8
kitchen-mixed-v0 1e-6
kitchen-partial-v0 1e-7
antmaze-medium-play-v2 1e-4
antmaze-medium-diverse-v2 1e-2
antmaze-large-play-v2 1e-4
antmaze-large-diverse-v2 1e-2
maze2d-umaze-v1 1e-7
maze2d-medium-v1 1e-8
maze2d-large-v1 1e-6
In Table 2, we report the average normalized scores on MuJoCo and AntMaze tasks when using a multi-modal Gaussian prior. The corresponding hyperparameter $\alpha$ are listed in Table 8.
Table 8: $\alpha$ used for multi-modal Gaussian prior on the D4RL benchmark.
$\alpha$ multi-modal
walker2d-medium-v2 0.01
walker2d-medium-replay-v2 0.01
walker2d-medium-expert-v2 0.001
hopper-medium-v2 0.01
hopper-medium-replay-v2 0.0001
hopper-medium-expert-v2 0.01
halfcheetah-medium-v2 0.0001
halfcheetah-medium-replay-v2 0.001
halfcheetah-medium-expert-v2 0.1
antmaze-medium-play-v2 0.1
antmaze-medium-diverse-v2 0.1
antmaze-large-play-v2 0.1
antmaze-large-diverse-v2 1.0
Appendix G Extensive Results
Full results of Figure 7(b) can be found in Table 9.
Table 9: Comparison of Prior Guidance with and without backpropagation through the denoising process (denoted as PG w/ Bp and PG w/o Bp, respectively) on the D4RL benchmark. We compute the mean and the standard error over 5 random seeds.
PG w/ Bp PG w/o Bp
[…]
Average 85.8 84.6
[…]
Average 81.0 81.3
[…]
Average 84.0 83.4
[…]
Average 164.2 164.6
Full results of Table 2 can be found in Table 10.
Table 10: Normalized scores of PG using multi-modal Gaussian prior on MuJoCo and Antmaze tasks. We compute the mean and the standard error over 5 random seeds.
multi-modal
[…]
Average 82.3
[…]
Average 85.8
Appendix H Ablation Stuides
We conducted ablation studies on the behavior regularization coefficient $\alpha$ of PG across MuJoCo, Kitchen, AntMaze, and Maze2D tasks. The normalized scores of PG for different values of $\alpha$ are reported in Table 11, 12, 13 and 14.
Table 11: Ablation study of $\alpha$ in MuJoCo tasks. We compute the mean and the standard error over 5 random seeds.
$\alpha$ 0.001 0.01 0.1 1.0 10.0
[…]
Table 12: Ablation study of $\alpha$ in Kitchen tasks. We compute the mean and the standard error over 5 random seeds.
$\alpha$ 0.1 1.0 10.0
[…]
Table 13: Ablation study of $\alpha$ in Antmaze tasks. We compute the mean and the standard error over 5 random seeds.
$\alpha$ 0.1 1.0 5.0 10.0 50.0
[…]
Table 14: Ablation study of $\alpha$ in Maze2D tasks. We compute the mean and the standard error over 5 random seeds.
$\alpha$ 0.001 0.01 0.1 1.0 10.0
[…]
We also conducted ablation studies on various behavior regularization functions $f$, as shown in Table 15.
Table 15: Comparison of normalized scores of Prior Guidance across four behavior regularization functions—KL, Reverse KL, and Pearson $\chi^{2}$—on the D4RL benchmark. We compute the mean and the standard error over 5 random seeds.
$f$ KL Reverse KL Pearson-$\mathcal{X}^{2}$
[…]
Average 84.6 85.0 82.9
[…]
Average 81.3 79.3 73.4
[…]
Average 83.4 82.7 78.1
[…]
Average 164.6 157.7 165.3
Appendix I Why PG underperform on standard MuJoCo tasks?
Table 16: Performance comparison across standard MuJoCo tasks.
DQL DV* PG (planner) PG (policy)
walker2d-medium-v2 87.0 79.5 82.3 84.5 $\pm$ 0.26
walker2d-medium-replay-v2 95.5 83.5 83.7 96.4 $\pm$ 2.49
walker2d-medium-expert-v2 110.1 109.0 109.4 109.8 $\pm$ 0.23
hopper-medium-v2 90.5 84.1 97.5 98.4 $\pm$ 1.01
hopper-medium-replay-v2 101.3 91.3 91.3 100.5 $\pm$ 1.27
hopper-medium-expert-v2 111.1 109.9 110.4 111.4 $\pm$ 0.57
halfcheetah-medium-v2 51.1 50.9 45.6 54.7 $\pm$ 0.42
halfcheetah-medium-replay-v2 47.8 46.4 46.4 48.6 $\pm$ 0.34
halfcheetah-medium-expert-v2 96.8 92.3 95.2 95.5 $\pm$ 0.26
Average 87.9 83.0 84.6 88.9
PG is a diffusion planner-based method designed for long-horizon planning, where capturing temporally extended dependencies is essential. However, MuJoCo locomotion tasks involve short-horizon, dense-reward control focused on making agents run faster, which does not require such lookahead capabilities [19]. As a result, diffusion planner-based methods, including PG, often underperform compared to diffusion policy-based approaches in these settings, as shown in Table 1. While the additional structure and trajectory-level modeling of planners benefit challenging long-horizon domains such as AntMaze, Kitchen, or Maze2D, they provide little advantage and can even be slightly detrimental in simpler environments where policies can already achieve near-optimal performance.
To examine whether the observed performance gap in MuJoCo tasks from the use of the diffusion planner rather than a limitation of the PG framework itself, we conduct an additional experiment by applying PG to a diffusion policy. Specifically, we replaced the standard Gaussian prior of a behavior-cloned diffusion policy with a learnable, behavior-regularized prior, following the same training procedure used in PG.
Table LABEL:tab:pg_policy shows that replacing the diffusion planner in PG with a diffusion policy (PG policy) consistently improved performance across MuJoCo tasks, raising the average score from 84.6 (PG planner) to 88.9 and surpassing the DQL baseline score of 87.9. This demonstrates that PG remains effective even in short-horizon environments, suggesting that the lower performance of PG in MuJoCo tasks is due to the nature of trajectory-level planning rather than a limitation of the method itself. The results for PG (policy) are averaged over 3 random seeds and reported as the mean $\pm$ standard error.